OpenStudy (anonymous):

Variables on both sides. Solve each equation. -2-(6m-5) = -12 -3m

OpenStudy (kl0723):

-2+6m+5 = -12-3m next...move all m terms to the left side :) 6m+3m = -12+2-5 next...add 9m = -15 divide to isolate m: m = -15/9 m = -5/3